OpenStudy (anonymous):

y varies inversely with x, and x = 150 when y = 3. Find the inverse variation equation and use it to find the value of x when y = 90. Answer 1/5 9/5 5 5/9

OpenStudy (jhannybean):

y varies inversely with x : y = k/x x = 150, y = 3 : 3 = k/150 k = 3(150) = 450 find x when y = 90 y = k/x 90 = 450/x x = 90/450 = ?

OpenStudy (anonymous):

0.2

OpenStudy (jhannybean):

what is 0.2? It's 2/10. What is the simplified version of 2/10?

OpenStudy (anonymous):

1/5

OpenStudy (jhannybean):

good job :)

OpenStudy (anonymous):

thanx